CASE 4–20 Ethics and the Manager, Understanding the Impact of Percentage Completion on Profit—Weighted-Average Method [Course Objective B] Gary Stevens and Mary James are production managers in the Consumer Electronics Division of General Electronics Company, which has several dozen plants scattered in locations throughout the world. Mary manages the plant located in Des Moines, Iowa, while Gary manages the plant in El Segundo, California. Production managers are paid a salary and get an additional bonus equal to 5% of their base salary if the entire division meets or exceeds its target profits for the year. The bonus is determined in March after the company’s annual report has been prepared and issued to stockholders.

Shortly after the beginning of the New Year, Mary received a phone call from Gary that went like this:

Gary: How’s it going, Mary?

Mary: Fine, Gary. How’s it going with you?

Gary: Great! I just got the preliminary profit figures for the division for last year and we are within $200,000 of making the year’s target profits. All we have to do is pull a few strings, and we’ll be over the top!

Mary: What do you mean? Gary: Well, one thing that would be easy to change is your estimate of the percentage completion of your ending work in process inventories. Mary: I don’t know if I can do that, Gary. Those percentage completion figures are supplied by Tom

Winthrop, my lead supervisor, who I have always trusted to provide us with good estimates.

Besides, I have already sent the percentage completion figures to corporate headquarters. Gary: You can always tell them there was a mistake. Think about it, Mary. All of us managers are doing as much as we can to pull this bonus out of the hat. You may not want the bonus check, but the rest of us sure could use it.

The final processing department in Mary’s production facility began the year with no work in process inventories. During the year, 210,000 units were transferred in from the prior processing department and 200,000 units were completed and sold. Costs transferred in from the prior department totaled $39,375,000. No materials are added in the final processing department. A total of $20,807,500 of conversion cost was incurred in the final processing department during the year.

Required:

1.            Tom Winthrop estimated that the units in ending inventory in the final processing department were 30% complete with respect to the conversion costs of the final processing department. If this estimate of the percentage completion is used, what would be the Cost of Goods Sold for the year? (Note: Since all units completed were sold, the cost of goods transferred out = Cost of Goods Sold.)

2.            Gary is recommending that the completion percentage by adjusted by 15 percentage points in order to assist the team in making their bonus.

a. Calculate the cost of goods sold if the ending inventory is 15% complete in regard to conversion costs. Would net income increase or decrease if this option was chosen over the 30% completion percentage? How much is the increase?

b. Calculate the cost of goods sold if the ending inventory is 45% complete in regard to conversion costs. Would net income increase or decrease if this option was chosen over the 30% completion percentage? How much is the increase?

c. Based on your calculations, which percentage is Gary suggesting that Mary use for her ending inventory calculations.

3.            Do you think Mary James should go along with the request to alter estimates of the percentage completion? Why or why not?
Answer #1

Total ending WIP in the final processing department = 210000-200000 = 10000 units.

Units of WIP completed = 30% of 10000 = 3000

Total goods on which cost incurred by final processing department = 200000+3000 = 203000.

Per unit cost incurred by final processing department = 20807500/203000 = $102.5

Per unit cost incurred by prior processing department = 39375000/210000 = $187.5

Cost of goods sold = {(102.5+187.5)x200000} = $58000000

2 (a). If completion percentage is 15% , then units completed by final processing department = 200000+15% of 10000 = 201500.

Cost per unit incurred by final processing department = 20807500/201500 = $103.26

Cost of goods sold = (187.5+103.26)x200000 = $58152605.

Since cost of goods sold has increased compared to what we got if completion percentage would have been 30% by 58152605-58000000 = $152605.

Thus, net income will decrease by the same amount that is by $152605.

(B) . If percentage completion is 45% then total units completed by final processing department = 200000+45% of 10000 = 204500

Cost per unit incurred by final processing department = 20807500/204500 = $101.75

Cost of goods sold = (187.5+101.75)x200000 = $57849633.

Thus cost of goods sold will decrease by $58000000-57849633 = $150367.

Thus, as a result net income will increase by $150367.

(C). Since gart wants that product tion manager to get bonus. And bonus is based on 5% of target profit. Profit will come more only when cost of goods sold is less. And we can clearly see when completion percentage is 45%, cost of goods sold is coming out to be less and profit is coming out to be more.

Thus, Gary will be suggesting Mary to use 45% for ending inventory calculation.

3 . Mary should not go along with the request to alter estimates of the percentage completion. The reason being

A. It will be unethical for the production manager to use a wrong completion than what should have been as he is working on a very responsible designation . And such incorrect reporting will amount to breach of trust by Mary James with the company.

2. Use of such percentage of completion my Mary would not give true and fair view of both profitability of the company and the WIP and inventory of the company and would adversely affect the decision making of the company.